OpenStudy (anonymous):

*I can help you with Psych, English, and Philosophy in return! In the below system, solve for y in the first equation.2x + y = 115x - 2y = -1 a.) 2x + 11 b.) -2x + 11 c.) 2x - 11 d.) -2x - 11

OpenStudy (anonymous):

Please help me?

OpenStudy (anonymous):

is this the system: 2x + y = 11 5x - 2y = -1 ???

OpenStudy (anonymous):

Is the equation written correctly? I just got an answer that isn't an option

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

2x + y = 115x - 2y = -1

OpenStudy (anonymous):

It has to be one of those four answers

OpenStudy (anonymous):

so if you're only solving for y in the first equation, subtract 2x from both sides

OpenStudy (anonymous):

Check

OpenStudy (anonymous):

\(\large 2x+y=11 \) \(\large 2x\color {red}{-2x}+y=11\color {red}{-2x} \) \(\large y=-2x+11 \)

OpenStudy (anonymous):

I think you ought to multiply the top equation by two and add the the equations

OpenStudy (anonymous):

nevermind me

OpenStudy (anonymous):

I GOT B!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

now... you must answer what's the meaning of life?

OpenStudy (anonymous):

The meaning of life is to find your own meaning.

OpenStudy (anonymous):

What keeps you going forward in life? The answer to that question could be your meaning.

OpenStudy (anonymous):

Are you there for your family? Your friends? Your job? Yourself? God?

OpenStudy (anonymous):

Maybe the meaning of life is to find pure happiness.

OpenStudy (anonymous):

Maybe the meaning of life is to help others find their own personal meaning of life.

OpenStudy (anonymous):

Maybe you have no meaning in life. Maybe you're here to just live and go about your days.